NvidiaIntely (nvidiaintely):

Which equivalence factor should you use to convert from 106 grams of O2 to moles of O2? (1 mol O2/16.00 grams O2) (16.00 grams O2/1 mol O2) (1 mol O2/32.00 grams O2) (32.00 grams O2/1 mol O2)

Nnesha (nnesha):

it's good to write it as a fraction because we have to cancel out units \[\rm (\frac{ 106 g ~of ~O_2 }{ 1 }) \cdot \frac{??}{???}\] we have to convert grams to mole so would you write grams at the numerator or mole for 2nd fraction ?

Nnesha (nnesha):

1 mole = atomic mass

NvidiaIntely (nvidiaintely):

Um... mole

Nnesha (nnesha):

correct so 1 mole = how many grams of O_2 remember O_2 means 2 oxygen

NvidiaIntely (nvidiaintely):

31.9988 grams.

Nnesha (nnesha):

ok so how would you write fraction ??

NvidiaIntely (nvidiaintely):

32.00 grams O2/1 mol O2, option D. P.S. how do I write a fraction on openstudy?

Nnesha (nnesha):

you can use the equation tool \[\huge\rm \frac{ 106 ~g~of~O_2 }{ 1 } \cdot \frac{ 32 ~g ~of~O_2}{1 ~mole ~of~O_2}\] is this correct ?

NvidiaIntely (nvidiaintely):

Yes.

Nnesha (nnesha):

okay simplify that fracation. would that give you the mole ?

Nnesha (nnesha):

reread my previous comments

Nnesha (nnesha):

whenever u solve conversation questions whatever word is to the right of the `to` will be the numerator convert kg `to `g (g would be at the numerator ) convert L `to` mL (mL would be the numerator ) or think about as an algebraic expression \[\frac{ x }{ 1 } \cdot \frac{x}{y} = \frac{x^2}{y}\] is different th\[\frac{ x }{ 1 } \cdot \frac{y}{x}\] ^^converting x `to` y

NvidiaIntely (nvidiaintely):

How do I simplify that fraction?

Nnesha (nnesha):

it's multiplication so multiply numerator by numerator and denominator ignore of O_2

Nnesha (nnesha):

\[\frac{ 106 g }{ 1 } \cdot \frac{32g}{1 mole }\] solve this would that give you the moles ??

NvidiaIntely (nvidiaintely):

\[\frac{ 3392 }{ 1 }\]?? I think it would give you the moles @Nnesha.

Nnesha (nnesha):

how would you multiply this x times x = ??

Nnesha (nnesha):

don't forget the units those are important.

Nnesha (nnesha):

whenever u solve conversation questions whatever word is to the right of the ` to` should be at the numerator convert kg to g (g would be at the numerator ) convert L to mL (mL would be the numerator ) here are some example

Nnesha (nnesha):

how would you multiply this x times x = ??

Nnesha (nnesha):

let me know when you are free it will be great idea if you finish this one first and then move on because the question u just posted is similar to this one .

NvidiaIntely (nvidiaintely):

I am ready now.

NvidiaIntely (nvidiaintely):

@Nnesha, I would like to understand this :)

Nnesha (nnesha):

okay i asked some questions.